OpenStudy (anonymous):

How many grams of oxygen gas will react completely with a block of calcium metal that is 3.0 cm by 3.5 cm by 4.2 cm, if the density of calcium is 1.55 g/mL? Show all steps of your calculation as well as the final answer. Ca + O2 → CaO

OpenStudy (anonymous):

Ca + O2 → CaO 1. The equation is NOT balanced as is. The balanced equation is 2Ca + O2 --------> CaO 2. Now, let's calculate the moles of Ca in the block...but first you have to get the mass of Ca. Volume = l * w* h = 3.0*3.5*4.2 = 44.1 cc or 44.1 mL 3. The density, you've been given already and it is 1.55 g/mL so to determine the gms of Ca you calculate mass = density*volume = 1.55 g/mL * 44.1 mL = 68.36 g of Ca. 4. Now, you have to calculate the mols of Ca and that is 68.36g/40.08g/mol = 1.71 mols of Ca. 5. The stoichiometry of the balanced equation states that you need 2 mols Ca for each mol O2 (a ratio of 2:1) but you've only got 1.71 mols of Ca to start with so you'll only react 1.71/2 mols of O2 and that is = 0.85 mol O2 and multiplying be O2 MW of 16.0 g/mol you'll get 13.6 g of O2.